# Fluctuations of the free energy of the spherical Sherrington–Kirkpatrick model with heavy-tailed interaction

### v1

@@ 3.2 Integral representation formula for the partition function
Since $G^{\prime}$ is an incresing function on the interval $(\lambda_{1},\infty)$ with
@@
Set $\gamma\in(\lambda_{1},\infty)$ be the (unique) solution to the equation $G^{\prime}(z)=0$. Then,
@@ 4 High temperature regime
In this section, we consider the regime $\lambda_{1}<b_{N}/2\beta$. Recall that we let $\gamma=\frac{b_{N}}{2\beta}(1+\frac{X_{N}}{N}+O(N^{-1-\epsilon}))$ in Lemma 3.17.
@@
For event $F_{1}=\left\{\lambda_{1}\leq\frac{b_{N}}{2\beta}\right\}$,
@@
Due to Lemma 3.5, we have $\lambda_{1}=\max\{|M_{ij}|\}(1+O(b_{N}^{-1/8})),\lambda_{N}=-\max\{|M_{ij}|\}(1+O(b_{N}^{-1/8}))$ holds with high probability. Hence, for $F_{1}^{\prime}=\{|M_{ij}|\leq\frac{b_{N}}{2\beta}\ \textit{for all }i,j\}$, $\mathbb{P}(F_{1}^{\prime}\Delta F_{1})\leq N^{-\delta}$ for some $\delta>0$. Also, for $F_{1}^{\prime\prime}=\{|M_{ij}|\leq\frac{b_{N}}{2\beta}(1-N^{-\epsilon})\textit{for all }i,j\}$ for small $\epsilon>0$, then $\mathbb{P}(F_{1}^{\prime\prime}\Delta F_{1})\leq N^{-\delta}$ for some $\delta>0.$ Therefore, we condition the matrix and all the statistics to event $F_{1}^{\prime\prime}$ in this section. For the convenience, we express variables same as before. Similar to the low temperature case, we first prove the following lemma for the derivatives of the function $G.$
@@
With high probability, the k-th derivative of G satisfies
@@
where $|\gamma-a|=O(b_{N}N^{-\delta})$ for some $\delta>0$.
@@
with high probability and Lemma 3.6 implies $\lambda_{1}<\frac{b_{N}}{2\beta}(1-N^{-\epsilon})$ holds for every $\epsilon>0$ from $\lambda_{1}<\frac{b_{N}}{2\beta}$. Since $\gamma=\frac{b_{N}}{2\beta}(1+\frac{X_{N}}{N})$ from Lemma 3.17, we conclude that
@@
Recall that the free energy is
@@
To handle the last term, we consider the following representation with a sufficiently small $\epsilon>0,$
$\int_{-\infty}^{+\infty}$ $\exp\left({\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}\right)dt=\sqrt{\frac{b_{N}}{N}}\int_{-\infty}^{+\infty}\exp\left({\frac{N}{2b_{N}}\left(G\left(\gamma+i\sqrt{\frac{b_{N}}{N}}t\right)-G(\gamma)\right)}\right)\mathrm{d}t$
@@
From Lemma 4.2 and Taylor approximation,
(4.1)$&=\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}\exp\left({-\frac{1}{4}G^{\prime\prime}(\gamma)t^{2}+O(\frac{N}{2b_{N}}\frac{1}{3!}\sup_{|a-\gamma|<b_{N}N^{\epsilon-1/2}}|G^{(3)}(a)|(\sqrt{b_{N}}t/\sqrt{N})^{3}}\right)\mathrm{d}t\\ &=\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}\exp\left({-\frac{1}{4}G^{\prime\prime}(\gamma)t^{2}+\frac{N}{2b_{N}}O(b_{N}^{-2}(b_{N}N^{\epsilon-1/2})^{3})}\right)\mathrm{d}t\\ &=\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}\exp\left({-\frac{1}{4}G^{\prime\prime}(\gamma)t^{2}+O(N^{-\epsilon})}\right)\mathrm{d}t\\ &=\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}e^{(-\frac{1}{4}G^{\prime\prime}(\gamma))t^{2}}\mathrm{d}t(1+O(N^{-\epsilon})).\end{split}$ (4.3)
Since $G^{\prime\prime}(\gamma)=O(b_{N}^{-1})$ and
@@
For the bound of (4.2), we have
|(4.2)$|&=\left|\int_{t>\sqrt{b_{N}}N^{\epsilon}}2\operatorname{\mathrm{Re}}\exp\left({\frac{N}{2b_{N}}(2\beta ti\sqrt{\frac{b_{N}}{N}}-\frac{b_{N}}{N}\sum_{i=1}^{N}\log(1-\frac{\sqrt{\frac{b_{N}}{N}}ti}{\gamma-\lambda_{i}}))}\right)\mathrm{d}t\right|\\ &\leq\int_{t>\sqrt{b_{N}}N^{\epsilon}}2\left|\exp\left({\frac{N}{2b_{N}}(2\beta ti\sqrt{\frac{b_{N}}{N}}-\frac{b_{N}}{N}\sum_{i=1}^{N}\log(1-\frac{\sqrt{\frac{b_{N}}{N}}ti}{\gamma-\lambda_{i}}))}\right)\right|\mathrm{d}t\\ &=\int_{t>\sqrt{b_{N}}N^{\epsilon}}2\exp\left({-\frac{1}{4}\sum_{i=1}^{N}\log(1+\frac{{\frac{b_{N}}{N}}t^{2}}{(\gamma-\lambda_{i})^{2}})}\right)\mathrm{d}t.\end{split}$ (4.4)
Note that there are at least $N/2$ eigenvalues that are less than equal to $b_{N}$ with high probability, since $\sum\lambda_{i}^{2}<b_{N}^{2}N^{\epsilon}$ with high probability. Using this fact, we further have
@@
$\leq\int_{\sqrt{b_{N}}N^{\epsilon}}^{b_{N}N}\exp\left({-\frac{b_{N}N^{2\epsilon}}{8N}}\right)\mathrm{d}t+\int_{t>b_{N}N}2\left(\frac{t^{2}}{b_{N}N}\right)^{-N/8}\mathrm{d}t$
$=O(b_{N}Ne^{-N^{2\epsilon}})+\left.\frac{1}{1-N/4}t^{-N/4+1}(c^{2}b_{N}N)^{N/8}\right|_{t=b_{N}N}^{t=\infty}$
$=O(b_{N}Ne^{-N^{2\epsilon}})+O(N^{-N/8}).$
Hence,
@@
since Stirling’s formula implies
@@
and it converges in probability to 00, which can be checked by an elementary calculation
$\lim_{N\to\infty}\left(\frac{x}{2}-\frac{N}{2}\log(1+\frac{x}{N})=0\right)$
for all $x>0$. Furthermore, in Lemma 4.3, we prove that $-\frac{1}{2}\sum_{i=1}^{N}\log(1-\lambda_{i}/\gamma)$ converges in distribution to some random variable $T$. In conclusion, if we define $F_{N}:=\log Z_{N}$, then it converges in distribution to a random variable $T$. ∎
@@
We conclude this section by proving that $T_{N}=-\sum_{i=1}^{N}\log(1-\lambda_{i}/\gamma)$ converges in distribution.
@@
The log statistics $T_{N}=-\sum_{i=1}^{N}\log(1-\lambda_{i}/\gamma)$ converges in distribution to some random variable.
@@
with high probability, and hence $|\lambda_{n}|,\lambda_{1}<\gamma$ with high probability. Thus,
@@
$|T_{N}|\leq\sum_{k=3}^{\infty}\frac{1}{k}|\frac{\Gamma}{\gamma}|^{k-2}S_{2}\leq-(\frac{\gamma}{\Gamma})^{2}{S_{2}}\log(1-\Gamma/\gamma).$
Since $S_{2}$ and ${\Gamma}/{\gamma}$ converge in distribution to some random variables, this proves $T_{N}$ converges to a non-degenerate random variable. ∎
@@ 5 Low temperature regime
In this section, we assume the event $F_{2}=\left\{\lambda_{1}>\frac{b_{N}}{2\beta}\right\}$ holds. Our goal in this section is to prove the following proposition:
@@
$\mathbb{P}(X>u)=\exp((2\beta)^{\alpha}-u^{-\alpha}),.$
@@
With high probability, the Taylor coefficient $G$ satisfies
$G^{(k)}(\gamma)=\Theta(N^{k-1}b_{N}^{-(k-1)}).$
@@
Recall that we proved $(\gamma-\lambda_{1})=\Theta(b_{N}/N)$ in Lemma 3.17. Then,
$G^{(k)}(\gamma)=\frac{b_{N}}{N}\sum_{i}\frac{(k-1)!}{(\gamma-\lambda_{i})^{k}}\geq\frac{b_{N}}{N}\frac{1}{(\gamma-\lambda_{1})^{k}}=\Theta(N^{k-1}b_{N}^{-k+1}).$
Since $\lambda_{1}-\lambda_{2}>b_{N}N^{-\epsilon}$ with high probability due to Lemma 3.6,
@@
We thus conclude that
$G^{(k)}(\gamma)=\Theta(N^{k-1}b_{N}^{-k+1}).$
We now consider the logarithm of the partition function,
@@
We claim that the last term in the equation above satisfies
@@
To prove this, we will show
@@
for some $C_{1},C_{2}>0$ and $c_{1},c_{2}\in\mathbb{R}$. For the upper bound, we have
@@
The remaining part is to show its lower bound. To show its lower bound, we will use contour integral. Consider the curve $\Gamma$ which satisfies $\operatorname{\mathrm{Im}}(G(t))=0$ on that curve and starting from $\gamma$ and end with $\infty$- ∞. By the Cauchy-Riemann equation $\operatorname{\mathrm{Re}}(G(z))$ is monotone along the curve. Since $G(z)$ decreases near $\gamma$ as imaginary part increases from its Taylor expansion, $G(z)=\operatorname{\mathrm{Re}}(G(z))$ decays along the curve.
We consider the path $\Gamma^{+}=\Gamma\cap\mathbb{C}^{+}$ and $\Gamma^{-}=\Gamma\cap\mathbb{C}^{-}$ which are starting from $\gamma$ and goes to $\infty$- ∞. Let $C_{R}=\{|z|=R\}\cap C^{+}\cap\operatorname{\mathrm{Re}}(z)<\gamma$, then $\operatorname{\mathrm{Re}}(G(z))\leq 2\beta\gamma-b_{N}\log(R/2)$
$|\int_{C_{R}}e^{\frac{N}{2b_{N}}G(z)}\mathrm{d}z|\leq\frac{T_{N}}{R^{N/2}-1}$
for some constant $T_{N}$ which only depends on $N$.
This implies
@@
since $\overline{G(\bar{z})}=G(z)$.
@@
To obtain the lower bound of K, we will prove the lower bound for the latter real part. Its Taylor series holds for $|t|<\frac{b_{N}}{N}$, and
$G(\gamma+t)=G(\gamma)+\sum_{k\geq 2}\frac{1}{k!}G^{(k)}(\gamma)t^{k}.$
On the curve $\Gamma^{+}$
@@
This implies below for $|t|\leq\frac{1}{N^{2}}$;
@@
$\leq O\left(\frac{N^{2}}{b_{N}^{2}}\right)(x^{2}+y^{2})|t|\leq O((\frac{N}{b_{N}})^{2}\frac{1}{N^{2}})(x^{2}+y^{2}).$
Set $\alpha=-\frac{G^{\prime\prime\prime}(\gamma)}{G^{\prime\prime}(\gamma)}>0$ and $\alpha=O(\frac{N}{b_{N}})$, then we have
00 0 $=x+\frac{1}{6}\alpha y^{2}-\frac{1}{2}\alpha x^{2}+O((\frac{N}{b_{N}})^{2}\frac{1}{N^{2}})(x^{2}+y^{2})$
$=x+\frac{1}{6}\alpha y^{2}(1+o(1))-\frac{1}{2}\alpha x^{2}(1+o(1))=x(1+o(1))+\frac{1}{6}\alpha y^{2}(1+o(1)).$
Therefore,
$x=-\frac{1}{6}\alpha y^{2}(1+o(1))\text{ for }|x+iy|\leq\frac{1}{N^{2}}$
and $\operatorname{\mathrm{Im}}y$ increases on $|t|<\frac{1}{N^{2}}$.
Since
$t^{2}=x^{2}+y^{2}=\{O(\frac{N}{b_{N}})y^{2}\}^{2}+y^{2}=O(\frac{N^{2}}{b_{N}^{2}})y^{4}+y^{2},$
we have
$y=\Theta(t),x=\Theta\left(\frac{N}{b_{N}}t^{2}\right)\text{ for }t<\frac{1}{N^{2}}.$
Now we will prove its lower bound using Lemma 5.3: for any real valued function f on $\Gamma^{+}$ which decreases along the curve as $z$ moves from $\gamma$ to $\infty$- ∞, $\int_{\Gamma^{+}}e^{f(z)}\mathrm{d}y\geq 0$. Hence, we construct below decreasing function and we show our integration is greater than their difference.
Let

### v2

@@ 3.2 Integral representation formula for the partition function
Since $G^{\prime}$ is an increasing function on the interval $(\lambda_{1},\infty)$ with
@@
Set $\gamma\in(\lambda_{1},\infty)$ as the (unique) solution to equation $G^{\prime}(z)=0$. Then,
@@ 4 High temperature regime
In this section, we consider the regime $\lambda_{1}<b_{N}/2\beta$. Recall that we let $\gamma=\frac{b_{N}}{2\beta}(1+\frac{X_{N}}{N})$ in Lemma 3.17.
@@
For event $F_{1}=\left\{\lambda_{1}<\frac{b_{N}}{2\beta}\right\}$,
@@
Throughout this section, we will assume that $|M_{ij}|<\frac{b_{N}}{2\beta}(1-N^{-\epsilon})$ for every $i,j$. To check the validity of this assumption, we first notice from the third part of Lemma 3.6 that for any fixed $\epsilon>0$, we have $\lambda_{1}\notin(\frac{b_{N}}{2\beta}(1-N^{-\epsilon})),\frac{b_{N}}{2\beta}(1+N^{-\epsilon}))$ with high probability. Since $\lambda_{1}<\frac{b_{N}}{2\beta}$, it implies that $\lambda_{1}<\frac{b_{N}}{2\beta}(1-N^{-\epsilon})$ with high probability. We also note from Lemma 3.5 that $\lambda_{1}=\max\{|M_{ij}|\}(1+O(b_{N}^{-1/8}))$ and $\lambda_{N}=-\max\{|M_{ij}|\}(1+O(b_{N}^{-1/8}))$ with high probability. Thus, given $F_{1}$, we find that $\max\{|M_{ij}|\}<\frac{b_{N}}{2\beta}(1-N^{-\epsilon})$ with high probability.
Our first result in this section is a lemma that provides an approximation for the $k$-th derivative of $G$.
@@
Fix $\delta>0$. For any $|\gamma-a|=O(b_{N}N^{-\delta})$, the $k$-th derivative of $G$ satisfies
@@
with high probability.
@@
with high probability. Moreover, if we assume that $\lambda_{1}<\frac{b_{N}}{2\beta}$, then from Lemma 3.6 we find that for any fixed $\epsilon>0$, $\lambda_{1}<\frac{b_{N}}{2\beta}(1-N^{-\epsilon})$ with high probability. Since $\gamma=\frac{b_{N}}{2\beta}(1+\frac{X_{N}}{N})$ from Lemma 3.17, we conclude that
@@
Recall from the integral representation formula for the partition function, Lemma 3.16,
@@
We first handle the last term of the equation above. For a sufficiently small $\epsilon>0$,
$\int_{-\infty}^{+\infty}$ $\exp\left({\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}\right)\mathrm{d}t=\sqrt{\frac{b_{N}}{N}}\int_{-\infty}^{+\infty}\exp\left({\frac{N}{2b_{N}}\left(G\left(\gamma+i\sqrt{\frac{b_{N}}{N}}t\right)-G(\gamma)\right)}\right)\mathrm{d}t$
@@
For (4.1), by considering the Taylor expansion of $G$ and applying Lemma 4.2,
$\begin{split}&\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}\exp\left({\frac{N}{2b_{N}}\left(G\left(\gamma+i\sqrt{\frac{b_{N}}{N}}t\right)-G(\gamma)\right)}\right)\mathrm{d}t\\ &=\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}\exp\left({-\frac{1}{4}G^{\prime\prime}(\gamma)t^{2}+O(\frac{N}{2b_{N}}\frac{1}{3!}\sup_{|a-\gamma|<b_{N}N^{\epsilon-1/2}}|G^{(3)}(a)|(\sqrt{b_{N}}t/\sqrt{N})^{3}}\right)\mathrm{d}t\\ &=\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}\exp\left({-\frac{1}{4}G^{\prime\prime}(\gamma)t^{2}+\frac{N}{2b_{N}}O(b_{N}^{-2}(b_{N}N^{\epsilon-1/2})^{3})}\right)\mathrm{d}t\\ &=\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}\exp\left({-\frac{1}{4}G^{\prime\prime}(\gamma)t^{2}+O(N^{-\epsilon})}\right)\mathrm{d}t\\ &=\int_{|t|\leq\sqrt{b_{N}}N^{\epsilon}}e^{(-\frac{1}{4}G^{\prime\prime}(\gamma))t^{2}}\mathrm{d}t(1+O(N^{-\epsilon})).\end{split}$ (4.3)
Since $G^{\prime\prime}(\gamma)=O(b_{N}^{-1})$, by applying the identity
@@
To estimate (4.2), we consider
$\left|\int_{|t|>\sqrt{b_{N}}N^{\epsilon}}\exp\left({\frac{N}{2b_{N}}\left(G\left(\gamma+i\sqrt{\frac{b_{N}}{N}}t\right)-G(\gamma)\right)}\right)\mathrm{d}t\right|$
$=\left|\int_{t>\sqrt{b_{N}}N^{\epsilon}}2\operatorname{\mathrm{Re}}\exp\left({\frac{N}{2b_{N}}(2\beta ti\sqrt{\frac{b_{N}}{N}}-\frac{b_{N}}{N}\sum_{i=1}^{N}\log(1-\frac{\sqrt{\frac{b_{N}}{N}}ti}{\gamma-\lambda_{i}}))}\right)\mathrm{d}t\right|$
$\leq\int_{t>\sqrt{b_{N}}N^{\epsilon}}2\left|\exp\left({\frac{N}{2b_{N}}(2\beta ti\sqrt{\frac{b_{N}}{N}}-\frac{b_{N}}{N}\sum_{i=1}^{N}\log(1-\frac{\sqrt{\frac{b_{N}}{N}}ti}{\gamma-\lambda_{i}}))}\right)\right|\mathrm{d}t$
$=\int_{t>\sqrt{b_{N}}N^{\epsilon}}2\exp\left({-\frac{1}{4}\sum_{i=1}^{N}\log(1+\frac{{\frac{b_{N}}{N}}t^{2}}{(\gamma-\lambda_{i})^{2}})}\right)\mathrm{d}t.$ (4.4)
Note that there are at least $N/2$ eigenvalues less than $\frac{b_{N}}{10\beta}$ with high probability, since $\sum\lambda_{i}^{2}<b_{N}^{2}N^{\epsilon}$ with high probability. Using this fact, we further have
@@
$\leq\int_{\sqrt{b_{N}}N^{\epsilon}}^{b_{N}N}\exp\left({-\frac{b_{N}C^{\prime}N^{2\epsilon}}{8N}}\right)\mathrm{d}t+\int_{t>b_{N}N}2\left(\frac{Ct^{2}}{b_{N}N}\right)^{-N/8}\mathrm{d}t$
$=O(b_{N}Ne^{-c^{\prime}N^{2\epsilon}})+\left.\frac{1}{1-N/4}t^{-N/4+1}(c^{2}b_{N}N)^{N/8}\right|_{t=b_{N}N}^{t=\infty}$
$=O(b_{N}Ne^{-c^{\prime}N^{2\epsilon}})+O(N^{-N/8})$
for some $N$-independent constants $C,C^{\prime},c,c^{\prime}$. Here, to obtain the inequality in the second line, we used $\log(1+x)\geq x/2$ for $x<1$ for the first part and $\exp(-a\log(1+b))\leq b^{-a}$ for $a,b>0$ for the second part.
Collecting the inequalities we obtained so far, we find
@@
where we used Stirling’s formula
@@
and it converges in probability to 00, which can be checked by an elementary fact
$\lim_{N\to\infty}\left(\frac{x}{2}-\frac{N}{2}\log(1+\frac{x}{N})\right)=0$
that holds for all $x>0$. In Lemma 4.3, we will prove that $-\frac{1}{2}\sum_{i=1}^{N}\log(1-\lambda_{i}/\gamma)$ converges in distribution to some random variable $T$. This shows that $F_{N}:=\log Z_{N}$ converges in distribution to the random variable $T$. Finally, $\lim_{N\to\infty}\mathbb{P}(F_{1})$ can be easily checked from Proposition 3.2. This concludes the proof of Proposition 4.1. ∎
@@
We conclude this section by proving that $T_{N}:=-\sum_{i=1}^{N}\log(1-\lambda_{i}/\gamma)$ converges in distribution.
@@
The log statistics $T_{N}=-\sum_{i=1}^{N}\log(1-\lambda_{i}/\gamma)$ converges in distribution to some ($N$-independent) random variable.
@@
with high probability, and hence $|\lambda_{n}|<\gamma$ and $\lambda_{1}<\gamma$ with high probability. Thus, with high probability, the following expansion holds:
@@
$|T_{N}|\leq\sum_{k=3}^{\infty}\frac{1}{k}|\frac{\Gamma}{\gamma}|^{k-2}S_{2}\leq-(\frac{\gamma}{\Gamma})^{2}{S_{2}}\log(1-\Gamma/\gamma)=:L_{N}.$
The random variable $L_{N}$ defined above converges to a non-degenerate random variable, since $S_{2}$ and ${\Gamma}/{\gamma}$ converge in distribution to some random variables.
We now rewrite $T_{N}$ as
$T_{N}=\sum_{i=1}^{N}-\log\left(1-\frac{\lambda_{i}}{\gamma}\right)-\frac{\lambda_{i}}{\gamma}-\frac{\lambda_{i}^{2}}{2\gamma^{2}}+S_{1}+\frac{1}{2}S_{2}$
and define the random variables
$U_{N}$ $=\sum_{i=1}^{N}-\log\left(1-\frac{\lambda_{i}}{\gamma}\right)-\frac{\lambda_{i}}{\gamma}-\frac{\lambda_{i}^{2}}{2\gamma^{2}},$
$U_{N}^{\epsilon}$ $=\sum_{|\lambda_{i}|>\epsilon\gamma}-\log\left(1-\frac{\lambda_{i}}{\gamma}\right)-\frac{\lambda_{i}}{\gamma}-\frac{\lambda_{i}^{2}}{2\gamma^{2}},$
$U_{N,k}=\sum_{i=1}^{k}\left(-\log\left(1-\frac{\lambda_{i}}{\gamma}\right)-\log\left(1-\frac{\lambda_{N+1-i}}{\gamma}\right)-\frac{\lambda_{i}+\lambda_{N+1-i}}{\gamma}-\frac{\lambda_{i}^{2}+\lambda_{N+1-i}^{2}}{2\gamma^{2}}\right).$
First, we notice that the difference between $U_{N}$ and $U_{N}^{\epsilon}$ is bounded by
$|U_{N}-U_{N}^{\epsilon}|$ $=\left|\sum_{|\lambda_{i}|<\epsilon\gamma}-\log\left(1-\frac{\lambda_{i}}{\gamma}\right)-\frac{\lambda_{i}}{\gamma}-\frac{\lambda_{i}^{2}}{2\gamma^{2}}\right|$
$\leq\sum_{|\lambda_{i}|<\epsilon\gamma}2\left|\frac{\lambda_{i}^{2}}{\gamma^{3}}\right|<2\epsilon\sum\frac{\lambda_{i}^{2}}{\gamma^{2}}=2\epsilon S_{2}.$
By definition, $U_{N,i}\leq U_{N,j}$ for $i\leq j$. Further, for any fixed $k$, if we let $M_{k}$ be the $k$-th large entry of $\{|M_{ij}|:1\leq i\leq j\leq N\}$, then with high probability, $\lambda_{k}=M_{k}(1+o(1))$ and $\lambda_{N+1-k}=-M_{k}(1+o(1))$. (See, e.g., [33].) We then find that $U_{N,k}$ converges in distribution to some random variable $V_{k}$. From the monotonicity of $U_{N,k}$, we find that $V_{1}\leq V_{2}\leq V_{3}\leq\ldots$. Thus, since $U_{N,k}\leq L_{N}$, from the convergence of $L_{N}$ we have
$\lim_{N\to\infty}V_{N}=V_{\infty}$
for some random variable $V_{\infty}$.
Let $F_{\epsilon}=\left\{\epsilon<S_{2}<\frac{1}{\sqrt{\epsilon}}\right\}$, then
$U_{N,1/\epsilon}\leq\left(U_{N}^{\epsilon}|F_{\epsilon}\right)\leq U_{N,\epsilon^{-5/2}}\text{ and }|U_{N}|F_{\epsilon}-U_{N}^{\epsilon}|F_{\epsilon}|\leq 2\epsilon S_{2}=2\sqrt{\epsilon},$
which also implies
$U_{N,1/\epsilon}-2\sqrt{\epsilon}\leq U_{N}|F_{\epsilon}\leq U_{N,\epsilon^{-5/2}}+2\sqrt{\epsilon}.$
Thus, we can obtain
$\mathbb{P}(U_{N,1/\epsilon}\geq U+2\sqrt{\epsilon})\leq\mathbb{P}(U_{N}|F_{\epsilon}\geq U)\leq\mathbb{P}(U_{N,\epsilon^{-5/2}}\geq U-2\sqrt{\epsilon})$
and this implies
$\lim_{\epsilon\to 0}\lim_{N\to\infty}\mathbb{P}(U_{N,1/\epsilon}\geq U+2\sqrt{\epsilon})=\lim_{\epsilon\to 0}\mathbb{P}(V_{1/\epsilon}\geq U+2\sqrt{\epsilon})=\mathbb{P}(V_{\infty}\geq U)$
$\lim_{\epsilon\to 0}\lim_{N\to\infty}\mathbb{P}(U_{N,\epsilon^{-5/2}}\geq U-2\sqrt{\epsilon})=\lim_{\epsilon\to 0}\mathbb{P}(V_{\epsilon^{-5/2}}\geq U-2\sqrt{\epsilon})=\mathbb{P}(V_{\infty}\geq U).$
Hence, we have
$\mathbb{P}(V_{\infty}\geq U)=\lim_{\epsilon\to 0}\mathbb{P}(V|F_{\epsilon}\geq U)=\lim_{\epsilon\to 0}\lim_{N\to\infty}P(U_{N}|F_{\epsilon}\geq U)=\mathbb{P}(V_{\infty}\geq U)$
We thus conclude that
$U_{N}\Rightarrow V_{\infty}.$
This proves that $T_{N}$ converges to a non-degenerate random variable. ∎
@@ 5 Low temperature regime
In this section, we assume that the event $F_{2}=\left\{\lambda_{1}>\frac{b_{N}}{2\beta}\right\}$ holds. Our goal in this section is to prove the following proposition:
@@
$\mathbb{P}(X>u)=\frac{1-\exp(-u^{-\alpha})}{1-\exp(-(2\beta)^{\alpha})}.$
@@
Recall the definition of $\gamma$ in Lemma 3.17. With high probability, there exist ($N$-independent) coefficients $C_{1}$ and $C_{2}$ such that the Taylor coefficients $G^{(k)}(\gamma)$ of $G$ at $\gamma$ satisfies
$C_{1}N^{k-1}b_{N}^{-(k-1)}N^{-k\epsilon}\leq G^{(k)}(\gamma)\leq C_{2}N^{k-1}b_{N}^{-(k-1)}.$
@@
Recall we proved in Lemma 3.17 that
$\gamma-\lambda_{1}=\frac{1}{2\beta-b_{N}/\lambda_{1}}\frac{b_{N}}{N}+O(b_{N}N^{-1-\epsilon}).$
We invoke Lemma 3.6 as in the proof of Lemma 4.2 to find that $\lambda_{1}>\frac{b_{N}}{2\beta}(1+N^{-\epsilon})$ with high probability. Thus, we have
$\gamma-\lambda_{1}=O(b_{N}N^{-1+\epsilon}).$
Further, since $2\beta-b_{N}/\lambda_{1}\leq 2\beta$, we have
$\gamma-\lambda_{1}\geq\frac{b_{N}}{2\beta N}+O(\frac{b_{N}}{N}N^{-\epsilon})\geq\frac{b_{N}}{4\beta N}.$
From the definition of $G^{(k)}(\gamma)$, we find an upper bound for $G^{(k)}(\gamma)$,
$G^{(k)}(\gamma)=\frac{b_{N}}{N}\sum_{i}\frac{(k-1)!}{(\gamma-\lambda_{i})^{k}}\geq\frac{b_{N}}{N}\frac{1}{(\gamma-\lambda_{1})^{k}}=\Theta(N^{k-1}b_{N}^{-k+1}N^{-k\epsilon}).$
For the lower bound for $G^{(k)}(\gamma)$, since $\lambda_{1}-\lambda_{2}>b_{N}N^{-\epsilon}$ with high probability due to Lemma 3.6,
@@
This completes the proof of the desired lemma. ∎
In order to prove Proposition 5.1, as in the high-temperature regime, we consider the integration representation formula,
@@
The key part of the proof is to show that
@@
To show this, we will prove the following claim:
@@
for some $C_{1},C_{2}>0$ and $c_{1},c_{2}\in\mathbb{R}$. For the upper bound in the claim, we notice
@@
To prove the lower bound of the claim, we will deform the integral by considering the following curve: let $\Gamma$ be the curve that starts from $\gamma$ and ends at $\infty$- ∞ such that for any point $t\in\Gamma$, we have $\operatorname{\mathrm{Im}}(G(t))=0$. (Implicit function theorem ensures the existence of such a curve.) By the Cauchy-Riemann equation, $\operatorname{\mathrm{Re}}(G(z))$ is monotone along the curve. Since $G(z)$ decreases near $\gamma$ as imaginary part increases from its Taylor expansion, $G(z)=\operatorname{\mathrm{Re}}(G(z))$ decays along the curve. We further decompose $\Gamma$ by $\Gamma=\Gamma^{+}\cup\Gamma^{-}$, where $\Gamma^{+}:=\Gamma\cap\mathbb{C}^{+}$ and $\Gamma^{-}:=\Gamma\cap\mathbb{C}^{-}$. Note that both $\Gamma^{+}$ and $\Gamma^{-}$ start from $\gamma$ and end at $\infty$- ∞. Let $C_{R}$ be an arc starting from $\{|z|=R\}\cap\Gamma^{+}$ to $\{|z|=R\}\cap\{\operatorname{\mathrm{Re}}(z)=\gamma\}$. Then for any sufficiently large $R$ with $R>2\lambda_{1}$, we have
$\operatorname{\mathrm{Re}}(G(z))=\operatorname{\mathrm{Re}}(2\beta z-\frac{b_{N}}{N}\sum_{i}\log(z-\lambda_{i}))\leq 2\beta\gamma-b_{N}\log(R/2).$
Then, the contour integral along the curve $C_{R}$ satisfies
$|\int_{C_{R}}e^{\frac{N}{2b_{N}}G(z)}\mathrm{d}z|\leq\frac{e^{N\beta\gamma/b_{N}}\pi R}{(R/2)^{N/2}}.$
In particular, as $R\to\infty$ the integral above converges to 00. This enables us to change the integral along $\operatorname{\mathrm{Re}}(z)=\gamma$ to the integral along the curve $\Gamma$, i.e.,
@@
where we also used the fact $\overline{G(\bar{z})}=G(z)$.
@@
To obtain a lower bound for $K$, we will prove the lower bound for the second term in the right side of the equation above. To estimate this term, we need to prove some properties of the contour $\Gamma^{+}$. We notice that the Taylor expansion
$G(\gamma+z)=G(\gamma)+\sum_{k\geq 2}\frac{1}{k!}G^{(k)}(\gamma)z^{k}$
holds for $|z|<\frac{b_{N}}{N}$. Thus, for $z=x+iy\in\Gamma^{+}$,
@@
We then find for $|z|\leq\frac{1}{N^{2}}$ that
@@
$\leq O\left(\frac{N^{2+2\epsilon}}{b_{N}^{2}}\right)(x^{2}+y^{2})|z|\leq O(\frac{N^{2\epsilon}}{b_{N}^{2}})(x^{2}+y^{2}).$
Set $a=-\frac{G^{\prime\prime\prime}(\gamma)}{G^{\prime\prime}(\gamma)}>0$. From Lemma 5.2, we find $a=O(\frac{N}{b_{N}}N^{2\epsilon})$ and then
00 0 $=x+\frac{1}{6}ay^{2}-\frac{1}{2}ax^{2}+O(\frac{N^{2\epsilon}}{b_{N}^{2}})(x^{2}+y^{2})$
$=x+\frac{1}{6}ay^{2}(1+o(1))-\frac{1}{2}ax^{2}(1+o(1))=x(1+o(1))+\frac{1}{6}ay^{2}(1+o(1)).$
Thus,
$x=-\frac{1}{6}ay^{2}(1+o(1))\text{ for }|x+iy|\leq\frac{1}{N^{2}}$
and $y$ increases along contour $\Gamma^{+}$ starting from 00 in $|z|<\frac{1}{N^{2}}$. Moreover, since
$|z|^{2}=x^{2}+y^{2}=y^{2}+O(\frac{N^{2+4\epsilon}}{b_{N}^{2}}y^{4}),$
with high probability, we have
$y=\Theta(|z|),x=\Theta\left(\frac{G^{\prime\prime\prime}(\gamma)}{G^{\prime\prime}(\gamma)}|z|^{2}\right)\text{ for }|z|<\frac{1}{N^{2}}.$
We now go back to the lower bound for
$2\int_{\Gamma^{+}}\exp(\frac{N}{2b_{N}}[G(z)-G(\gamma)])\mathrm{d}y.$
We decompose $\Gamma^{+}$ to $\Gamma_{1}\cup\Gamma_{2}\cup\Gamma_{3}$, where we let